OpenStudy (anonymous):

write an equation in slope-intercept form for the line that passes through(-1,-2) and (3,-7)

OpenStudy (anonymous):

(-9/2)I believe

OpenStudy (blurbendy):

first find the slope: -7 + 2/ 3 + 1 = -5/4 then use the point-slope form: y - y1 = slope(x - x1) y +2 = -5/4(x +1) y + 2 = -5/4x -5/4 y = -5/4x -13/4

OpenStudy (blurbendy):

and now it's in slope-intercept form

OpenStudy (anonymous):

thank you

OpenStudy (blurbendy):

yw